# Numerical efficiency of explicit time integrators for phase-field models

### v1

@@ 1 Introduction
We note that the time evolution of different processes in the microstructure may span several orders of magnitude, making the simple forward Euler scheme often quite inefficient. This may simply stem different timescales of the equations, but may also couple into externally given control of a field parameter. Examples of the latter include e.g. the thermal control during solidification and sintering, where process times may be on the order of minutes to hours depending on the cooling/heating rate. Resolving the microstructure with relevant diffusivities then often puts the stable time step on the order of microseconds or less, making full-process simulations expensive to unfeasible. Another consideration, especially for sintering, is that the process speed depends on the grain size $G$ as a power function. For example, if densification occurs via grain boundary diffusion the densification speed can scale as $G^{-4}$Rahaman [2017]. Hence increasing the scale of simulation from e.g. a grid spacing of 1 to 10 makes this process $10^{4}$ times slower, but a forward Euler scheme gains only a factor of $10^{2}$ in its stable timestep, which means simulating to equivalent density takes 100 times as long! As the physical process occurs that much slower, unconditionally stable integrators can easily increase their time step substantially with little precision loss.
For the present paper, the focus will be on models with so-called obstacle potentials, which take on finite energy values in part of the domain, but infinite outside of it. This causes the evolution of the phase-field to be trivial outside of a narrow-band called the interface, with the remaining part of the domain being called bulk. Identification of the narrow-band can then be used to speed up simulations by not requiring the calculation of updates outside of it. The time integration needs to account for the infinite potential, which is often done using a projection approachGarcke et al. [1999], Daubner et al. [2023].
Although there is a large literature of integrators specifically crafted for the Cahn-Hilliard equationGomez and Hughes [2011], Christlieb et al. [2014], Vignal et al. [2017], Xu et al. [2019], Guillén-González and Tierra [2014], Cheng et al. [2021], Botchev et al. [2024] and many more, where the fourth-order nature makes the stiffness much more acute, a thorough efficiency comparison between them is more rare Christlieb et al. [2014], Vignal et al. [2017], Cheng et al. [2021], Botchev et al. [2024]; the focus is generally on assuring energy stability, i.e. that the monotonic decrease of energy predicted by the continuous problem also holds in the discrete case. This applies even more so for the Allen-Cahn equationVignal et al. [2017], Xu et al. [2019], Cheng et al. [2021], with no comparative efficiency investigations being found in the literature at all for obstacle potentials with explicit integrators. Hence this work intends to start to fill this gap by answering the following set of questions:
@@
To answer these questions, benchmark geometries with analytical solutions are defined, which allow for a more appropriate performance comparison than using a reference solution. However, these do suffer from usually being restricted to simpler geometry and parameter combinations. Hence a final geometry, grain growth with and without mobile or immobile particles, is also considered as a more complex geometry spanning several orders of magnitudes in its parameters.
@@ 2 Model
A phase-field model using an obstacle potential will be employed within the present work, with coupling to mass diffusion realized via a Kim-Kim-Suzuki approachKim et al. [1999]. The independent variables are the phase-field $\phi(x,t)$ with $N$ entries and the concentration $c(x,t)$ with $K=1$ independent entries whose arguments will be dropped for brevity. The energy functional for this reads
$F$ $=\int a+w+fdV$ (1)
$a$ $=-\sum_{\alpha>\beta}A_{\alpha\beta}\nabla\phi_{\alpha}\nabla\phi_{\beta}$ (2)
$w$ $=\begin{cases}\sum_{\alpha>\beta}B_{\alpha\beta}\phi_{\alpha}\phi_{\beta}&\phi\in\Delta^{N}\\ \infty&else\end{cases}$ (3)
$f$ $=\sum h_{\alpha}(\phi)g_{\alpha}(c_{\alpha})$ (4)
@@
$\Delta^{N}=\{(\phi_{0},\ldots,\phi_{N-1}):\sum\phi_{\alpha}=1\wedge\phi_{\alpha}\geq 0,\forall\alpha\in\{0,\ldots,N-1\}\}$ (5)
@@
The time evolution of $\phi$ is assumed to follow a variational derivative subject to the constraint $\sum\phi_{\alpha}=1$. This constraint is built into the time evolution following Steinbach and Pezzolla [1999] by a pairwise construction
$\frac{\partial\phi_{\alpha}}{\partial t}=\frac{1}{N_{z}}\sum_{\beta}L_{\alpha\beta}(\frac{\delta{F}}{\delta{\phi_{\alpha}}}-\frac{\delta{F}}{\delta{\phi_{\beta}}})$ (6)
with the pairwise mobility $L_{\alpha\beta}$ and the number of phases $N_{z}$ which have nonzero gradients at a point. Reducing the above equation for two phases while accounting for the constraint and demanding equilibrium, one obtains an analytically solvable ODE with a solution of the form
@@
In order to allow for an arbitrary number $N$ of phase-fields to be simulated efficiently, at each point in the domain only a fixed number of phase-fields $N_{p}$ is stored, in the spirit of Kim et al. [2006]. This is allowable due to the finite extent of the phase-field, which takes on trivial values on a large part of the domain. Thus each phase-field cell has $N_{p}$ phase-fields and integers indicating the phase these belong to, as well as an integer $N_{z}\leq N_{p}$ counting phase-fields with non-zero gradients. The number $N_{p}$ is adjusted to the problem at hand, with the benchmarks comparing to analytical solutions also being cheaply computable without this scheme since there $N\leq 3$. For these cases this storage scheme is unnecessary, but was still used for consistency with the final benchmark of section 4.5 where $N_{p}=12$ is employed. The resulting runtimes for benchmarks will thus be larger than if a simple direct $N\in\{2,3\}$ scheme were to be used; the number of right-hand side evaluations however is unaffected by this. Note that even though $N_{p}$ phases are formally saved, only the local phases $N_{z}$ are loaded and used for computation. This is in contrast to remapping schemes such as Permann et al. [2016] which need to load and calculate the entire chosen length, even if their update is trivially zero or vanishingly small.
Due to infinite energy outside of the simplex in the energy eq. 3 any $\phi$ that is calculated has to be within the simplex. This is accomplished by the following trivial projection: For any $\phi_{\alpha}<0$, set its value to zero; if any $\phi_{\alpha}\geq 1$, set it to one and set all other $\phi_{\beta}$ to zero. After this part the $\sum\phi_{\alpha}=1$ constraint may no longer be fulfilled, which is accounted for by dividing the $\phi$ vector by its sum. For the final geometry described in section 4.5 we employ a mobility-based simplex as proposed in Daubner et al. [2023]. A more detailed investigation of the influence of the chosen simplex together with mobility contrasts is left to future work.
@@
$=\nabla\cdot M_{c}\nabla\sum_{\alpha}h(\phi_{\alpha},\phi)\frac{\partial g_{\alpha}}{\partial c}$ (9)
with some mobility $M_{c}$. As $c$ is the independent variable whereas the Gibbs energies are formulated in their phase concentrations, some additional relation is required. For this the concentration is additionally represented as
$c$ $=\sum h_{\alpha}c_{\alpha}$ (10)
@@
i.e. $c$ is a weighted sum of the phase concentrations and local equilibrium is assumed. We may thus simplify
$\sum_{\alpha}h(\phi_{\alpha},\phi)\frac{\partial g_{\alpha}}{\partial c}=\mu$ (13)
@@
$=\nabla\cdot\sum_{\alpha}(D_{\alpha}h_{\alpha}(\frac{\partial^{2}g_{\alpha}}{\partial c_{\alpha}^{2}})^{-1}\nabla\mu)$ (15)
$=\nabla\cdot\sum_{\alpha}(D_{\alpha}h_{\alpha}\nabla c_{\alpha})$ (16)
in which the form for $M_{c}=\sum_{\alpha}D_{\alpha}h_{\alpha}(\frac{\partial^{2}g_{\alpha}}{\partial c_{\alpha}^{2}})^{-1}$ is chosen such that it cancels the relevant part from $\nabla\mu=\frac{\partial^{2}g_{\alpha}}{\partial c_{\alpha}^{2}}\nabla c_{\alpha}$. The last line writes out the sum and then replaces $\nabla\mu$ with its equivalent in terms of $\nabla c_{\alpha}$ before moving back into sum form. One can thus interpret the concentration evolution as a weighted average of phase-diffusion problems. The phase-wise concentrations $c_{\alpha}$ may be obtained from solving the system described by eqs. 10 and 11.
@@ 3 Methods
In this section the time integrators used to solve eqs. 6 and 16 are introduced together with some additional implementation details. The numerical framework used for the investigation supports execution on both the CPU and GPU, with parallelization achieved using the Message Passing Interface (MPI). It will be detailed in another publication. Spatial discretization is generally carried out with second order finite differences, with divergences being resolved as a sum of flux differences between faces. Point values are laid out on a cell-centered regular Cartesian grid with spacing $\Delta x$, with each unknown in a cell representing a degree of freedom (DoF).
@@ 3.1 Temporal discretization
A variety of explicit temporal integrators is evaluated in the present paper, with a short overview given in the present section. Among them is the “standard” forward Euler integrator as used by many explicit phase-field codes as a reference to how much faster other integrators may be; integrators suited for integration together with advective terms as their stability region includes a part of the imaginary axis (SSP schemes); and finally unconditionally stable explicit integrators built for parabolic problems (STS schemes). STS schemes are chosen as they allow for more efficient time integration for semi-stiff problems and SSP schemes for potential inclusion of advective terms, which will be the focus of a later work. SSP schemes may also be seen representative of “normal” Runge-Kutta methods which improve temporal accuracy but do not gain computational efficiency (no. RHS evaluations vs. stability limit) over the forward Euler integrator.
@@
The simplest integrator considered is the forward Euler (FE, FEuler) integrator given by
@@
in which the stability polynomial $R(\Delta t\lambda)$ is introduced. Explicit Runge-Kutta (RK) methods may always be written in the form eq. 22, with the only difference being the expression for $R$. In order for a method to be absolutely stable, i.e. errors introduced earlier do not grow exponentially, the boundLeVeque [2007]
@@
We also consider two strong positivity preserving (SSP) schemes, namely SSP(n)2 and SSP(10)4, which are respectively an n-stage scheme of order 2 and a 10-stage scheme of order 4. These schemes are efficient and accurate integrators for advective terms. The implementation is similar to the one given in Ketcheson [2008] except that an explicit additional vector is used since fulfilling the storage assumption is quite difficult on GPUs. Following the cited work, we have the stability limits
@@
Finally, super timestepping (STS) schemes of order 1 and 2 are also considered within the work. Either scheme with $s$ stages may be written in the general form
@@
with schemes of order 1 having $\nu_{j}=0,\tilde{\gamma_{j}}=0$. The coefficients $\mu_{j},\nu_{j},\tilde{\mu_{j}},\tilde{\gamma_{j}}$ depend on the chosen stability polynomial basis and the order of the scheme. A common feature however is that the stable time step scales as
@@
i.e. quadratically in the number of stages $s$ relative to the stable FE step (largest eigenvalue). Thus these schemes can be said to be unconditionally stable, as one might simply choose a large number of stages to achieve the desired stability limit. Choosing a larger step also makes the scheme more efficient relative to the FE scheme as increasingly fewer RHS evaluations will be required; this is in contrast to the SSP(n)2 scheme, whose stability domain only grows linearly. Within this work we only consider the Lagrange basis introduced by Meyer et al. [2012, 2014] for simplicity; the general trends should follow for other bases, with a possible exception of a Chebyshev basis when using spatially variable coefficients, as these may introduce step artifacts in the solutionMeyer et al. [2014]. In order to estimate the stable Euler time step $\Delta t_{e}$ necessary to compute the stable number of stages via eq. 31, we employ Gershgorin’s circle theorem on eqs. 6 and 16 separately for a two-phase system. This results in estimates $\lambda_{\phi},\lambda_{c}$ which are detailed in the supplementary material. The larger of the two resulting eigenvalues is then used to estimate a stable Euler step $t_{e}=\frac{2}{\lambda}$, which is multiplied by a safety factor of $0.9$ before being used to estimate the stage number, which is rounded to the next largest odd number as even stage numbers fail at damping short wavelengthsMeyer et al. [2012]. If a problem only involves the phase-field, as in section 4.3, only the phase-field eigenvalue estimate is employed without terms relating to the chemical driving force. For simplicity no nonlinear power method was used for estimating the eigenvalues; for more complicated couplings in which Gershgorin’s circle theorem yields very crude bounds or involves substantial linearization it may also speed up the computation by allowing larger steps.

### v2

@@ 1 Introduction
We note that the time evolution of different processes in the microstructure may span several orders of magnitude, making the simple forward Euler scheme often quite inefficient. This may simply stem different timescales of the equations, but may also couple into externally given control of a field parameter. Examples of the latter include e.g. the thermal control during solidification and sintering, where process times may be on the order of minutes to hours depending on the cooling/heating rate. Resolving the microstructure with relevant diffusivities then often puts the stable time step on the order of microseconds or less, making full-process simulations expensive to unfeasible. Another consideration, especially for sintering, is that the process speed depends on the grain size $G$ as a power function. For example, if densification occurs via grain boundary diffusion the densification speed can scale as $G^{-4}$Rahaman2017 . Hence increasing the scale of simulation from e.g. a grid spacing of 1 to 10 makes this process $10^{4}$ times slower, but a forward Euler scheme gains only a factor of $10^{2}$ in its stable timestep, which means simulating to equivalent density takes 100 times as long! As the physical process occurs that much slower, unconditionally stable integrators can easily increase their time step substantially with little precision loss.
For the present paper, the focus will be on models with so-called obstacle potentials, which take on finite energy values in part of the domain, but infinite outside of it. This causes the evolution of the phase-field to be trivial outside of a narrow-band called the interface, with the remaining part of the domain being called bulk. Identification of the narrow-band can then be used to speed up simulations by not requiring the calculation of updates outside of it. The time integration needs to account for the infinite potential, which is often done using a projection approachGarcke1999 , Daubner2023 .
Although there is a large literature of integrators specifically crafted for the Cahn-Hilliard equationGomez2011 , Christlieb2014 , Vignal2017 , Xu2019 , Guillen-Gonzalez2014 , Cheng2021 , Botchev2024 and many more, where the fourth-order nature makes the stiffness much more acute, a thorough efficiency comparison between them is more rare Christlieb2014 , Vignal2017 , Cheng2021 , Botchev2024 ; the focus is generally on assuring energy stability, i.e. that the monotonic decrease of energy predicted by the continuous problem also holds in the discrete case. This applies even more so for the Allen-Cahn equationVignal2017 , Xu2019 , Cheng2021 , with no comparative efficiency investigations being found in the literature at all for obstacle potentials with explicit integrators. Hence this work intends to start to fill this gap by answering the following set of questions:
@@
To answer these questions for particular cases, benchmark geometries with analytical solutions are defined, which allow for a more appropriate performance comparison than using a reference solution. In contrast to the benchmarks suggested by Jokisaari2017 the analytical solutions allow for exact and cheap calculation of the error with respect to the sharp interface problem, which allows the differentiation of space, time and interface width errors and is independent of the chosen free energy formulation. Furthermore, these can be seen as extensions for benchmarking Allen-Cahn problems with optional coupling to diffusion equations without requiring 4th order Cahn-Hilliard equations to be solved.
However, these do suffer from usually being restricted to simpler geometry and parameter combinations. Hence a final geometry, grain growth with and without mobile or immobile particles, is also considered as a more complex geometry spanning several orders of magnitudes in its parameters.
@@ 2 Model
A phase-field model using an obstacle potential will be employed within the present work, with coupling to mass diffusion realized via a Kim-Kim-Suzuki approachKim1999 . The independent variables are the phase-field $\phi(x,t)$ with $N$ entries and the concentration $c(x,t)$ with $K=1$ independent entries whose arguments will be dropped for brevity. The energy functional for this reads
$F$ $=\int f_{v}dV=\int a+w+fdV$ (1)
$a$ $=-\sum_{\alpha=0}^{N-1}\sum_{\alpha>\beta}^{N-1}A_{\alpha\beta}\nabla\phi_{\alpha}\nabla\phi_{\beta}$ (2)
$w$ $=\begin{cases}\sum_{\alpha}^{N-1}\sum_{\alpha>\beta}^{N-1}B_{\alpha\beta}\phi_{\alpha}\phi_{\beta}&\phi\in\Delta^{N}\\ \infty&else\end{cases}$ (3)
$f$ $=\sum_{\alpha=0}^{N-1}h_{\alpha}(\phi)g_{\alpha}(c_{\alpha})$ (4)
@@
$\Delta^{N}=\{(\phi_{0},\ldots,\phi_{N-1}):\sum_{\alpha=0}^{N-1}\phi_{\alpha}=1\wedge\phi_{\alpha}\geq 0,\forall\alpha\in\{0,\ldots,N-1\}\}$ (5)
@@
The time evolution of $\phi$ is assumed to follow a variational derivative subject to the constraint $\sum_{\alpha=0}^{N-1}\phi_{\alpha}=1$. This constraint is built into the time evolution following Steinbach1999 by a pairwise construction
$\frac{\partial\phi_{\alpha}}{\partial t}$ $=\frac{1}{N_{z}}\sum_{\beta=0}^{N-1}L_{\alpha\beta}n(\phi_{\alpha},\nabla\phi_{\alpha})n(\phi_{\beta},\nabla\phi_{\beta})(\frac{\delta{F}}{\delta{\phi_{\alpha}}}-\frac{\delta{F}}{\delta{\phi_{\beta}}})$ (6)
$\frac{\delta{F}}{\delta{\phi_{\alpha}}}$ $=\frac{\partial f_{v}}{\partial\phi_{\alpha}}-\nabla\cdot\frac{\partial f_{v}}{\partial\nabla\phi_{\alpha}}$ (7)
with the pairwise mobility $L_{\alpha\beta}$ and the number of phases $N_{z}$ which have nonzero gradients at a point.
This construction can be thought of as a constrained gradient flow on the manifold $\sum_{\alpha=0}^{N-1}\phi_{\alpha}=1$. Note that the sum here is also only taken over phases which have nonzero gradients, i.e. whose evolution is not trivial. To keep the sum formulation consistent with this, the function $n(\phi_{\alpha},\nabla\phi_{\alpha})$ is introduced which is 1 if this phase-field is nontrivial, but zero otherwise.
Reducing eq. 6 for two phases while accounting for the sum constraint and demanding equilibrium, one obtains an analytically solvable ODE with a solution of the form
@@
$=\nabla\cdot M_{c}\nabla\sum_{\alpha=0}^{N-1}h(\phi_{\alpha},\phi)\frac{\partial g_{\alpha}}{\partial c}$ (10)
with some mobility $M_{c}$. As $c$ is the independent variable whereas the Gibbs energies are formulated in their phase concentrations $c_{\alpha}$, some additional relation is required. For this the concentration is additionally represented as
$c$ $=\sum_{\alpha=0}^{N-1}h_{\alpha}c_{\alpha}$ (11)
@@
i.e. $c$ is a weighted sum of the phase concentrations and local equilibrium is assumed by equality of the chemical potential across phases. We may thus simplify
$\sum_{\alpha=0}^{N-1}h(\phi_{\alpha},\phi)\frac{\partial g_{\alpha}}{\partial c}=\mu$ (14)
@@
$=\nabla\cdot\sum_{\alpha=0}^{N-1}(D_{\alpha}h_{\alpha}(\frac{\partial^{2}g_{\alpha}}{\partial c_{\alpha}^{2}})^{-1}\nabla\mu)$ (16)
$=\nabla\cdot\sum_{\alpha=0}^{N-1}(D_{\alpha}h_{\alpha}\nabla c_{\alpha})$ (17)
in which the form for $M_{c}=\sum_{\alpha=0}^{N-1}D_{\alpha}h_{\alpha}(\frac{\partial^{2}g_{\alpha}}{\partial c_{\alpha}^{2}})^{-1}$ is chosen such that it cancels the relevant part from $\nabla\mu=\frac{\partial^{2}g_{\alpha}}{\partial c_{\alpha}^{2}}\nabla c_{\alpha}$. The last line writes out the sum and then replaces $\nabla\mu$ with its equivalent in terms of $\nabla c_{\alpha}$ before moving back into sum form. One can thus interpret the concentration evolution as a weighted average of phase-diffusion problems. The phase-wise concentrations $c_{\alpha}$ may be obtained from solving the system described by eqs. 11 and 12.
@@ 3 Methods
In this section the time integrators used to solve eqs. 6 and 17 are introduced together with some additional implementation details. The numerical framework used for the investigation supports execution on both the CPUs and GPUs, with parallelization achieved using the Message Passing Interface (MPI). It will be detailed in another publication. Spatial discretization is generally carried out with second order finite differences, with divergences being resolved as a sum of flux differences between faces. Point values are laid out on a cell-centered regular Cartesian grid with spacing $\Delta x$, with each unknown in a cell representing a degree of freedom (DoF).
3.1 Phase-field details
In order to allow for an arbitrary number $N$ of phase-fields to be simulated efficiently, at each point in the domain only a fixed number of phase-fields $N_{p}$ is stored, in the spirit of Kim2006 . This is allowable due to the finite extent of the phase-field, which takes on trivial values on a large part of the domain. Thus each phase-field cell has $N_{p}$ phase-fields and integers indicating the phase these belong to, as well as an integer $N_{z}\leq N_{p}$ counting phase-fields with non-zero gradients. The number $N_{p}$ is adjusted to the problem at hand, with the benchmarks comparing to analytical solutions also being cheaply computable without this scheme since there $N\leq 3$. For these cases this storage scheme is unnecessary, but was still used for consistency with the final benchmark of section 5.3 where $N_{p}=12$ is employed. The resulting runtimes for benchmarks will thus be larger than if a simple direct $N\in\{2,3\}$ scheme were to be used; the number of right-hand side evaluations however is unaffected by this. Note that even though $N_{p}$ phases are formally saved, only the local phases $N_{z}$ are loaded and used for computation. This is in contrast to remapping schemes such as Permann2016 which need to load and calculate the entire chosen length, even if their update is trivially zero or vanishingly small. We furthermore note that the filter $n(\phi_{\alpha},\nabla\phi_{\alpha})$ in eq. 6 is naturally adapted by this reduced storage scheme, as only nontrivial values are stored.
Due to infinite energy outside of the simplex in the energy eq. 3 any $\phi$ that is calculated has to be within the simplex. This is accomplished by the following trivial projection: For any $\phi_{\alpha}<0$, set its value to zero; if any $\phi_{\alpha}\geq 1$, set it to one and set all other $\phi_{\beta}$ to zero. After this part the $\sum_{\alpha=0}^{N-1}\phi_{\alpha}=1$ constraint may no longer be fulfilled, which is accounted for by dividing the $\phi$ vector by its sum. For the final geometry described in section 5.3 we employ a mobility-based simplex as proposed in Daubner2023 . A more detailed investigation of the influence of the chosen simplex together with mobility contrasts is left to future work.
@@ 3.2 Temporal discretization
A variety of explicit temporal integrators is evaluated in the present paper, with a short overview given in the present section. Among them is the “standard” forward Euler integrator as used by many explicit phase-field codes as a reference to how much faster other integrators may be.
Another integrator group is those suited for integration together with advective terms, as their stability region includes a part of the imaginary axis and they have the so-called strong stability property (SSP), the SSP schemes; due to this property, monotonic properties of the continuous time evolution can be preserved, which is critical for hyperbolic problemsKetcheson2008 . Finally, unconditionally stable integrators built for the solution of parabolic problems are also employed, the so-called super timestepping (STS) schemesSommeijer1998 , Meyer2012 , Meyer2014 .
STS schemes are chosen as they allow for more efficient time integration for semi-stiff problems and SSP schemes for potential inclusion of advective terms, which will be the focus of a later work. SSP schemes may also be seen representative of “normal” Runge-Kutta methods which improve temporal accuracy but do not gain computational efficiency (no. RHS evaluations vs. stability limit) over the forward Euler integrator.
@@
The simplest integrator considered is the forward Euler (FE, FEuler) integratorLeVeque2007 given by
@@
in which the stability polynomial $R(\Delta t\lambda)$ is introduced. Explicit Runge-Kutta (RK) methods may always be written in the form eq. 23, with the only difference being the expression for $R$. In order for a method to be absolutely stable, i.e. errors introduced earlier do not grow exponentially, the boundLeVeque2007
@@
We also consider two strong positivity preserving (SSP) schemes, namely SSP(n)2 and SSP(10)4, which are respectively an n-stage scheme of order 2 and a 10-stage scheme of order 4, both of which are described in Ketcheson2008 . These schemes are efficient and accurate integrators for advective terms. The implementation is effectively the one given in Ketcheson2008 except that an explicit additional vector is used since fulfilling the storage assumption is quite difficult on GPUs. For example, the SSP(n)2 scheme is written in pseudocode in algorithm 1 with $q_{3}$ being introduced as an additional buffer to evade the problem of reading and writing to $q_{1}$ concurrently. CPU-focused codes may avoid this by computing $\operatorname{rhs}(q_{1})$ in a layered fashion and writing back to $q_{1}$ as cells enter and exit the domain of dependence. Note that for the phase-field, the simplex constraint is applied to every stage of the solution and hence the stages evolve on the appropriate sub-manifold. The structure for the SSP(10)4 scheme is similar to the SSP(n)2 scheme and hence will not be repeated here.
$q_{1}=u^{n}$
$q_{2}=u^{n}$
$q_{3}$ // extra buffer
for $i=1$ to s1s-1 do
  $q_{3}=q_{1}+\Delta t\frac{\operatorname{rhs}(q_{1})}{s-1}$
  if $u$ is phase-field then
   simplex($q_{3}$)
  end if
  swap_buffers($q_{3}$, $q_{1}$)
end for
$q_{1}=\frac{(s-1)q_{3}+q_{2}+\Delta t\operatorname{rhs}(q_{3})}{s}$
$u^{n+1}=q_{1}$
Algorithm 1 Pseudocode for the SSP(s)2 scheme using an extra buffer variable.
Following Ketcheson2008 , we have the stability limits
@@
Finally, super timestepping (STS) schemes of order 1 and 2 are also considered within the work.
These are Runge-Kutta schemes whose stability polynomial $R$ as introduced earlier is constructed in such a way to allow the stability region to grow superlinearly with the number of stagesVerwer1990 , Meyer2012 . These schemes use orthogonal polynomials (e.g. Chebyshev polynomials Sommeijer1998 , Legendre polynomials Meyer2012 , Meyer2014 , …) as the stability polynomials $R_{j}$ of each stage $j$ in the resulting Runge-Kutta schemeVerwer1990 . By relying on the recursive properties of the chosen orthogonal polynomials, these schemes can be written in a recursive form, e.g. for $s$ stagesMeyer2014 :
@@
that is to say, the zeroth stage $u^{n,0}$ is the the old solution and the first stage $u^{n,1}$ is a scaled Euler step. The remaining stages form a sequence which depends on only a limited number of prior stages to compute their stage solutions $u^{n,j}$, which means that even though hundreds of stages may be used, the used memory is limited to a small number of solution vectors, which makes it a low-storage Runge-Kutta method. The final stage $s$ represents the solution at $t_{n+1}=t_{n}+\Delta t$. The coefficients $\mu_{j},\nu_{j},\tilde{\mu_{j}},\tilde{\gamma_{j}}$ depend on the chosen stability polynomial basis and the order of the scheme, with schemes of order 1 having $1-\mu_{j}-\nu_{j}=0,\tilde{\gamma_{j}}=0$ and hence no need to store information about the zeroth stage. An implementation of an STS1 scheme with $s$ stages is given in algorithm 2 which uses a periodic index $j$ over a fixed number of buffers $q$. The STS2 scheme is quite similar, except that the old solution and initial RHS are also stored and the update uses the full expression from eq. 31.
$q$ // 3 stage buffers
$q[0]=u^{n}$
$q[1]=q[0]+\Delta t\tilde{\mu_{j}}\operatorname{rhs}(q[0])$
for $i=2$ to $s$ do
  $j=i\mod 3;j_{p}=(i-1)\mod 3;j_{pp}=(i-2)\mod 3;$
  $q[j]=\mu_{j}q[j_{p}]+\nu_{j}q[j_{pp}]+\Delta t\tilde{\mu_{j}}\operatorname{rhs}(q[j_{p}])$
  if $u$ is phase-field then
   simplex($q[j]$)
  end if
end for
$u^{n+1}=q[j]$
Algorithm 2 Pseudocode for the STS1 scheme.
A common feature of STS schemes is that the stable time step scales as
@@
i.e. quadratically in the number of stages $s$ relative to the stable FE step (largest eigenvalue). Thus these schemes can be said to be unconditionally stable, as one might simply choose a large number of stages to achieve the desired stability limit. Choosing a larger step also makes the scheme more efficient relative to the FE scheme as increasingly fewer RHS evaluations will be required; this is in contrast to the SSP(n)2 scheme, whose stability domain only grows linearly. Within this work we only consider the Legendre basis introduced by Meyer2012 , Meyer2014 for simplicity, with schemes thus becoming Runge-Kutta-Legendre (RKL) schemes; the general trends should follow for other bases, with a possible exception of a Chebyshev basis when using spatially variable coefficients, as these may introduce step artifacts in the solutionMeyer2014 . In order to estimate the stable Euler time step $\Delta t_{e}$ necessary to compute the stable number of stages via eq. 32, we employ Gershgorin’s circle theorem on eqs. 6 and 17 separately for a two-phase system. This results in estimates $\lambda_{\phi},\lambda_{c}$ which are detailed in the supplementary material. The larger of the two resulting eigenvalues is then used to estimate a stable Euler step $t_{e}=\frac{2}{\lambda}$, which is multiplied by a safety factor of $0.9$ before being used to estimate the stage number, which is rounded to the next largest odd number as even stage numbers fail at damping short wavelengthsMeyer2012 . If a problem only involves the phase-field, as in section 5.2.1, only the phase-field eigenvalue estimate is employed without terms relating to the chemical driving force. For simplicity no nonlinear power method was used for estimating the eigenvalues; for more complicated couplings in which Gershgorin’s circle theorem yields very crude bounds or involves substantial linearization it may also speed up the computation by reducing the necessary stage count.